OpenStudy (anonymous):

Did I simplify this math problem correctly? (2a - 1) - (-5a - 5) Do I add it and have it be 7a + 4?

OpenStudy (anonymous):

-3a+4

OpenStudy (anonymous):

wait no im wrong ur right i was looking at ti wrong lol

OpenStudy (anonymous):

Anyone else?

OpenStudy (anonymous):

Yes. 2a-1 + (-1)(-5a) + (-1)(-5) = (2+5)a -1 + 5 = 7a+4